Mr. smith invested $24,000, part at 6% and the rest at 10%. if the interest earned was $2,000, find the amount invested at each rate.
Answer by checkley77(12844)

.10x+.06(24,000-x)=2,000
.10x+1,440-.06x=2,000
.04x=2,000-1,440
.04x=560
x=560/.04
x=14,000 invested @ 10%.
24,000-14,000=10,000 invested @ 6%.
Proof:
.10*14,000+.06*10,000=2,000
1,400+600=2,000
2,000=2,000